# Logarithmic de Rham stacks and non-abelian Hodge theory

### v1

@@ 3.1.1. The construction of $p$-curvature
where the first equivalence above is from Notation 2.44, and the second equivalence is Cartier duality.
@@
For computations, it is very helpful to know that $p$-curvature makes sense at the level of abelian categories.
Proposition 3.5.
[…]
Proof.
We defined $\psi_{p}$ as a composition of functors; each of these functors is $t$-exact, as
  1. (1)
[…]
  2. (2)
    the equivalence $\mathcal{D}_{\operatorname{qc}}(F_{X/S}^{*}(X/S)^{\widehat{\operatorname{dR}}})=\mathcal{D}_{\operatorname{qc}}(B\mathbb{V}(F_{X}^{*}\mathbb{L}_{X/S}^{\vee})^{\#,\wedge})$ is induced by an isomorphism of stacks, and hence it is $t$-exact;
  3. (3)
    the equivalence $\mathcal{D}_{\operatorname{qc}}(B\mathbb{V}(F_{X}^{*}\mathbb{L}_{X/S}^{\vee})^{\#,\wedge})=\mathcal{D}_{\operatorname{qc}}(\mathbb{V}(F_{X}^{*}\mathbb{L}_{X/S})$ is known to be $t$-exact for the torsion $t$-structure on the right and the standard $t$-structure on the left.
Hence the composition is $t$-exact. ∎
@@ 3.1.2. Treating $p$-curvature as a morphism
on $(X/S)^{\widehat{\operatorname{dR}}}$? And the answer is fortunately yes. TODO: FIX THIS FAITHFUL
@@ 3.2. Relative Cartier descent
As $\pi:X\to(X/S)^{\widehat{\operatorname{dR}}}$ is a flat cover, by flat descent we have that $\epsilon$ is an equivalence if and only if $\pi^{*}(\epsilon)$ is an equivalence. But $\pi^{*}(\epsilon)$ is exactly the morphism () 3.8, and so conditions 1 and 2 are equivalent.
@@
It follows from the commutativity of () 3.10 that
@@
Thus () 3.11 tells us $\pi^{*}(\psi_{p}(\nu^{*}\mathcal{F}))=0,$ which since $\pi$ is a flat cover implies $\psi_{p}(\nu^{*}\mathcal{F})=0.$ In particular, condition 1 implies condition 3.
@@ 4.1.4. The BNR correspondence
We now prove a variant of the Beauville-Narasimhan-Ramanan (BNR) correspondence for stacky curves, which we will use in our study of logarithmic non-abelian Hodge theory. Beauville-Narasimhan-Ramanan [4] proved the first version of the BNR correspondence for smooth spectral curves, and Schaub [32] later generalized their work to arbitrary special curves. Our treatment follows Schaub [32].
@@ 4.1.5. Functoriality of Hitchin morphisms
When comparing our logarithmic non-abelian Hodge theorem to de Catalado-Zhang’s [14], it will be useful to have the following result.
@@ 4.2. On the characteristic polynomial of $p$-curvature
Using de Rham stacks give us some advantage here; indeed, in the usual story of characteristic polynomials of $p$-curvature (see Laszlo-Pauly [24], and also Zhang [13]), one views the $p$-curvature as a $F^{*}\Omega^{1}_{X/S}$-twisted linear endomorphism on $X$; this makes it non-obvious that the Hitchin base for $p$-curvature is correct, and indeed one has to prove a theorem asserting some specialness property of $p$-curvature. In our de Rham stack approach, as we shall see, we need no theorems: just definitions.
@@ 5.1. Recollections on Olsson’s approach to logarithmic geometry
None of the results in this section are original, and all are essentially due to Olsson [31]; we recall them here merely for the reader’s convenience. In preparing this section, the author also found the works of Borne [11], [12], Esnault-Groechenig [19] and Laaroussi [23] very helpful.
@@ 5.1.1. Root stacks and Frobenius twists
The same lemma also appears as Lemma 3.16 of Esnault-Groechenig’s [19]; though we reproduce the short proof for the reader’s benefit.
@@ 5.4.1. On $(\operatorname{pt}/B\mathbb{G}_{m})^{\widehat{\operatorname{dR}}}$
The key point is that formation of both de Rham and sheared de Rham stacks preserve quasi-syntomic covers.
@@
In characteristic 0, finite type group schemes are smooth; in positive characteristic, finite type group schemes are quasi-syntomic. Thus $G\to\operatorname{Spec}(k)$ is a quasisynotomic surjection; hence $\operatorname{Spec}(k)\to BG$ is a quasisyntomic surjection. As formation of $\operatorname{dR}$ preserves quasisynotomic covers, we deduce that $\operatorname{Spec}(k)\to(BG)^{\operatorname{dR}}$ is still a quasi-syntomic cover. Hence we have a surjection $\operatorname{pt}\to(BG)^{\operatorname{dR}}$ whose fiber is $G^{\operatorname{dR}}$; this is exactly the data needed to identify $B(G^{\operatorname{dR}})\simeq(BG)^{\operatorname{dR}},$ so we conclude. The same argument works for $\widehat{\operatorname{dR}}$.
@@
Warning 5.25.
The sequence () 5.24 is not split as a short exact sequence of group schemes! We only construct a scheme-theoretic section; it will not respect the group laws. In fact, no section can be a group homomorphism, because the Cartier dual to () 5.24 is the Artin-Schreier sequence
@@
Sanath Devalapurkar, in a private communication to the author, explained that the non-sheared version of () 5.24 splits group theoretically, and hence $(\operatorname{pt}/B\mathbb{G}_{m})^{\operatorname{dR}}\to B\mu_{p}$ is the trivial gerbe. We will not need this in this paper, so we do not include it.
@@
The Artin-Schreier sequence appearing in Warning 5.25 allows us to compute $p$-curvature in terms of residue.
@@
As in Warning 5.25, after applying Cartier duality this inclusion becomes the Artin-Schreier cover; as $\psi_{p}$ and residue are both defined in terms of Cartier duality, we deduce that $\psi_{p}$ is the image of $\operatorname{res}$ under Artin-Schreier, and hence $\psi_{p}=\operatorname{res}^{p}-\operatorname{res},$ as desired. ∎
@@ 6.2.3. Intermezzo: a de Rham BNR correspondence
de Cataldo-Zhang ([14] Lemma 4.1) prove a very similar theorem, but without using these stacky spectral curves. The relationship between our two theorems is explained by the following remark: if $\mathcal{F}$ is a torsion free coherent sheaf on $C^{\prime}_{a}$ (the non-stacky spectral curve) with a $\mathcal{D}_{a}$-module structure, then there is a unique way to endow $\mathcal{F}$ with a parabolic structure compatible with the $\mathcal{D}_{a}$-module structure. Indeed, this is more or less the argument given to prove Lemma 6.19, which will be a crucial ingredient in our proof.
@@ 6.3. The more precise comparison
To conclude our proof of Theorem 6.29, we will essentially quotient the equivalence () 6.31 by $\mathcal{P}.$ This strategy is slightly different than Groechenig’s [20] original approach, and is essentially identical to the non-logarithmic argument made by de Cataldo-Groechenig-Zhang [15] in their Section 4.2.
@@
6.4. de Cataldo-Zhang’s Hitchin base
In the rest of this section, we will explain how to use our logarithmic non-abelian Hodge theorem to prove de Cataldo-Zhang’s [14] logarithmic non-abelian Hodge theorem.
@@
Definition 6.32 (de Cataldo-Zhang’s [14] Hitchin base).
Following de Cataldo-Zhang [14], we let $\mathcal{A}^{\operatorname{dCZ}}$ denote the fiber product
@@
where the map $\operatorname{AS}$ is just the map applying Artin-Schreier to each component.
@@
which we call de Cataldo-Zhang’s Hitchin morphism.
The stack of splittings $\mathcal{S}$ still appears in de Cataldo-Zhang’s logarithmic non-abelian Hodge theorem. However, they use a smaller Picard group.
@@
de Cataldo-Zhang also use a smaller Dolbeault side than we do.
@@
Finally, we can state de Cataldo-Zhang’s theorem.
@@
Identifying $\mathcal{S}\simeq\mathcal{M}_{\operatorname{dR}}^{\operatorname{reg}}$ as in Lemma 6.23 and using de Cataldo-Zhang’s Hitchin morphism $\mathcal{M}_{\operatorname{dR}}\xrightarrow{h_{\operatorname{dCZ}}}\mathcal{A}^{\operatorname{dCZ}},$ we can view $\mathcal{S}$ as a $\mathcal{A}^{\operatorname{dCZ}}$-stack.
Theorem 6.38 (de Cataldo-Zhang [14]).
@@
$\mathcal{S}\times^{\mathcal{P}^{\operatorname{dCZ}}}\tilde{\mathcal{M}_{\operatorname{Dol}}^{\operatorname{dCZ}}}\to\mathcal{M}_{\operatorname{dR}},$
@@ 6.5. Residues and parabolic structures
6.5.1. Deducing de Cataldo-Zhang’s theorem

### v2

@@ 3.1.1. The construction of $p$-curvature
where the first equivalence above is from Notation 2.47, and the second equivalence is Cartier duality.
@@ 3.1.2. Treating $p$-curvature as a morphism
on $(X/S)^{\widehat{\operatorname{dR}}}$? And the answer is fortunately yes.
@@ 3.2. Relative Cartier descent
As $\pi:X\to(X/S)^{\widehat{\operatorname{dR}}}$ is a flat cover, by flat descent we have that $\epsilon$ is an equivalence if and only if $\pi^{*}(\epsilon)$ is an equivalence. But $\pi^{*}(\epsilon)$ is exactly the morphism () 3.7, and so conditions 1 and 2 are equivalent.
@@
It follows from the commutativity of () 3.9 that
@@
Thus () 3.10 tells us $\pi^{*}(\psi_{p}(\nu^{*}\mathcal{F}))=0,$ which since $\pi$ is a flat cover implies $\psi_{p}(\nu^{*}\mathcal{F})=0.$ In particular, condition 1 implies condition 3.
@@ 4.1.4. The BNR correspondence
We now prove a variant of the Beauville–Narasimhan–Ramanan (BNR) correspondence for stacky curves, which we will use in our study of logarithmic non-abelian Hodge theory. Beauville–Narasimhan–Ramanan [4] proved the first version of the BNR correspondence for smooth spectral curves, and Schaub [37] later generalized their work to arbitrary special curves. Our treatment follows Schaub [37].
@@ 4.1.5. Functoriality of Hitchin morphisms
When comparing our logarithmic non-abelian Hodge theorem to de Catalado–Zhang’s [15], it will be useful to have the following result.
@@ 4.2. On the characteristic polynomial of $p$-curvature
Using de Rham stacks give us some advantage here; indeed, in the usual story of characteristic polynomials of $p$-curvature (see Laszlo-Pauly [25], and also Zhang [14]), one views the $p$-curvature as a $F^{*}\Omega^{1}_{X/S}$-twisted linear endomorphism on $X$; this makes it non-obvious that the Hitchin base for $p$-curvature is correct, and indeed one has to prove a theorem constraining the possible characteristic polynomials of $p$-curvature to get the Hitchin bases to align. In our de Rham stack approach, this constraint just comes from the fact that the $p$-curvature actually lives on $X^{\operatorname{dR}}.$
@@ 5.1. Recollections on Olsson’s approach to logarithmic geometry
None of the results in this section are original, and all are essentially due to Olsson [36]; we recall them here merely for the reader’s convenience. In preparing this section, the author also found the works of Borne [12], [13], Esnault–Groechenig [20] and Laaroussi [24] very helpful.
@@ 5.1.1. Root stacks and Frobenius twists
The same lemma also appears as Lemma 3.16 of Esnault–Groechenig’s [20]; though we reproduce the short proof for the reader’s benefit.
@@ 5.3. The log Azumaya property
For the reader’s benefit, we include an illustration of the main computation used in Hablicsek’s proof.
Example 5.14 (The Azumaya property for $\mathbb{A}^{1}$).
Let $S=\mathbb{A}^{1}_{k}/\mathbb{G}_{m,k}$ for $k$ an $\mathbb{F}_{p}$-algebra, and let $X=\operatorname{Spec}k[x].$ Take $D$ the divisor cut out by $x=0.$ The logarithmic Frobenius twist $X^{(1)}$ is the $p^{\text{th}}$ root stack of $X^{\prime}=\operatorname{Spec}k[x^{p}]$ along the divisor $x^{p}=0.$
Then $\mathcal{D}_{X}(\log D)=k[x,x\partial],$ with the two generators having commutator
$x\partial\cdot x-x\cdot x\partial=x.$
Note that $(x\partial)^{p}-x\partial$ and $x^{p}$ are both central elements of $\mathcal{D}_{X}(\log D).$
The pushforward $F_{X/S,*}\mathcal{D}_{X}(\log D),$ thought of as a quasicoherent sheaf on $X^{\prime}$ with parabolic structure (in the sense of Borne [13]), is just the $k[x^{p}]$-algebra $k[x,x\partial]$ with parabolic structure
$\cdots\subseteq F^{1}=x\cdot k[x,x\partial]\subseteq F^{0}=k[x,x\partial]\subseteq F^{-1}=x^{-1}\cdot k[x,x\partial]\subseteq F^{-2}=x^{-2}\cdot k[x,x\partial]\subseteq\cdots.$
The logarithmic cotangent bundle of $X^{(1)}$ is just the trivial bundle. Writing $\mathbb{V}(\omega_{X^{(1)}/S})=X^{(1)}\times_{k}\operatorname{Spec}k[\zeta],$ we can then upgrade $F_{X/S,*}\mathcal{D}_{X}(\log D)$ to an algebra on $\mathbb{V}(\omega_{X^{(1)}/S})$ by having $\zeta$ act on $F_{X/S,*}\mathcal{D}_{X}(\log D)$ as multiplication by the central element $(x\partial)^{p}-x\partial.$ In this way, we view $F_{X/S,*}\mathcal{D}_{X}(\log D)$ as a quasicoherent algebra on $\mathbb{V}(\omega_{X^{(1)}/S}).$
As the map $X\to X^{\prime}$ is faithfully flat, the induced map $X[\sqrt[p]{pD}]\to X^{(1)}$ of $p^{\text{th}}$ root stacks is faithfully flat as well. As being Azumaya can be checked on a flat cover, it suffices to prove that the pullback of $F_{X/S,*}\mathcal{D}_{X}(\log D)$ to $X[\sqrt[p]{pD}]\times\mathbb{A}^{1}_{\zeta}$ is a matrix algebra. As a sheaf on $X\times\mathbb{A}^{1}_{\zeta}$ with parabolic structure, this pullback $\mathcal{D}^{\prime}$ can be modelled as the parabolic sheaf
$F^{i}\mathcal{D}^{\prime}=\sum_{j}(k[x,x\partial]\cdot x^{j})\otimes_{k[x^{p},(x\partial)^{p}-x\partial]}(k[x,(x\partial)^{p}-x\partial]\cdot x^{i-j}),$
where $\zeta$ acts as $(x\partial)^{p}-x\partial.$
View $\mathcal{D}_{X}(\log D)$ as a sheaf on $X[\sqrt[p]{pD}]\times\mathbb{A}^{1}_{\zeta}$ via the parabolic bundle
$\cdots\subseteq F^{1}=k[x,x\partial]\cdot x\subseteq F^{0}=k[x,x\partial]\subseteq F^{-1}=k[x,x\partial]\cdot x^{-1}\subseteq\cdots,$
where again we have $\zeta$ acting as $(x\partial)^{p}-x\partial.$ The structure sheaf of $X[\sqrt[p]{pD}]\times\mathbb{A}^{1}_{\zeta}$ can be identified as the subbundle
$\cdots\subseteq F^{1}=k[x,(x\partial)^{p}-x\partial]\cdot x\subseteq F^{0}=k[x,(x\partial)^{p}-x\partial]\subseteq F^{-1}=k[x,(x\partial)^{p}-x\partial]\cdot x^{-1}\subseteq\cdots$
of $\mathcal{D}_{X}(\log D).$ We denote this subbundle by $\mathcal{A}_{X}.$
Consider now the sheaf of endomorphisms $\underline{\operatorname{End}}_{\mathcal{A}_{X}}(\mathcal{D}_{X}(\log D)),$ viewing $\mathcal{D}_{X}(\log D)$ as a right $\mathcal{A}_{X}$-module. There is a natural map of parabolic bundles
$\mathcal{D}^{\prime}\to\underline{\operatorname{End}}_{\mathcal{A}_{X}}(\mathcal{D}_{X}(\log D)),$
which sends $(f\cdot x^{j})\otimes(g\otimes x^{i-j})\in F^{i}\mathcal{D}^{\prime}$ to the endomorphism
$D\mapsto(fx^{j})D(gx^{i-j}).$
This endomorphism is $\mathcal{A}_{X}$-linear; indeed, it commutes with $D\mapsto Dx$ since $g\in k[x,(x\partial)^{p}-x\partial]$ commutes with $x,$ and it commutes with right multiplication by $(x\partial)^{p}-x\partial$ for the same reason.
This endomorphism lies in $F^{i}\underline{\operatorname{End}}_{\mathcal{A}_{X}}(\mathcal{D}_{X}(\log D))$ because if $D\in F^{n}\mathcal{D}_{X}(\log D),$ then $fx^{j}Dgx^{i-j}$ lies in $F^{n+i}\mathcal{D}_{X}(\log D)$ (as $x^{j},x^{i-j}$ change degree in a predictable way, and $f,g$ both act only by increasing degree). This map is an isomorphism by the same argument used by Bezuknivakov [5].
@@ 5.4.1. On $(\operatorname{pt}/B\mathbb{G}_{m})^{\widehat{\operatorname{dR}}}$
The key point is that formation of both de Rham and sheared de Rham stacks preserve quasisyntomic covers; see Lemma 6.3 of Bhatt–Lurie [9] for a slightly stronger claim in the non-sheared case. In the sheared case, a similar argument is used to show the analogous result in the upcoming work on sheared prismatization [8].
@@
In characteristic 0, finite type group schemes are smooth; in positive characteristic, finite type group schemes are quasisyntomic. Thus $G\to\operatorname{Spec}(k)$ is a quasisynotomic surjection; hence $\operatorname{Spec}(k)\to BG$ is a quasisyntomic surjection. As formation of $\operatorname{dR}$ preserves quasisynotomic covers, we deduce that $\operatorname{Spec}(k)\to(BG)^{\operatorname{dR}}$ is still a quasisyntomic cover. Hence we have a surjection $\operatorname{pt}\to(BG)^{\operatorname{dR}}$ whose fiber is $G^{\operatorname{dR}}$; this is exactly the data needed to identify $B(G^{\operatorname{dR}})\simeq(BG)^{\operatorname{dR}},$ so we conclude. The same argument works for $\widehat{\operatorname{dR}}$.
@@
Warning 5.26.
The sequence () 5.25 is not split as a short exact sequence of group schemes! We only construct a scheme-theoretic section; it will not respect the group laws. In fact, no section can be a group homomorphism, because the Cartier dual to () 5.25 is the Artin-Schreier sequence
@@
Sanath Devalapurkar, in a private communication to the author, explained that the non-sheared version of () 5.25 splits group theoretically, and hence $(\operatorname{pt}/B\mathbb{G}_{m})^{\operatorname{dR}}\to B\mu_{p}$ is the trivial gerbe. We will not need this in this paper, so we do not include it.
@@
The Artin-Schreier sequence appearing in Warning 5.26 allows us to compute $p$-curvature in terms of residue.
@@
As in Warning 5.26, after applying Cartier duality this inclusion becomes the Artin-Schreier cover; as $\psi_{p}$ and residue are both defined in terms of Cartier duality, we deduce that $\psi_{p}$ is the image of $\operatorname{res}$ under Artin-Schreier, and hence $\psi_{p}=\operatorname{res}^{p}-\operatorname{res},$ as desired. ∎
@@ 6.2.3. Intermezzo: a de Rham BNR correspondence
de Cataldo–Zhang ([15] Lemma 4.1) prove a very similar theorem, but without using these stacky spectral curves. The relationship between our two theorems is explained by the following remark: if $\mathcal{F}$ is a torsion free coherent sheaf on $C^{\prime}_{a}$ (the non-stacky spectral curve) with a $\mathcal{D}_{a}$-module structure, then there is a unique way to endow $\mathcal{F}$ with a parabolic structure compatible with the $\mathcal{D}_{a}$-module structure. Indeed, this is more or less the argument given to prove Lemma 6.19, which will be a crucial ingredient in our proof.
@@ 6.3. The more precise comparison
To conclude our proof of Theorem 6.29, we will essentially quotient the equivalence () 6.31 by $\mathcal{P}.$ This strategy is slightly different than Groechenig’s [21] original approach, and is essentially identical to the non-logarithmic argument made by de Cataldo–Groechenig–Zhang [16] in their Section 4.2.
@@
6.4. de Cataldo–Zhang’s Hitchin base
In the rest of this section, we will explain how to use our logarithmic non-abelian Hodge theorem to prove de Cataldo–Zhang’s [15] logarithmic non-abelian Hodge theorem.
@@
Definition 6.32 (de Cataldo–Zhang’s [15] Hitchin base).
Following de Cataldo–Zhang [15], we let $\mathcal{A}^{\operatorname{dCZ}}$ denote the fiber product
@@
where the map $\operatorname{AS}$ is obtained as follows: viewing an element of $\mathbb{A}^{r}$ as the coefficients of a degree $r$ monic polynomial $p$, let $\operatorname{AS}(p)$ be the coefficients of the polynomial whose roots are these Artin-Scrheier images of the roots of $p.$
@@
which we call de Cataldo–Zhang’s Hitchin morphism.
The stack of splittings $\mathcal{S}$ still appears in de Cataldo–Zhang’s logarithmic non-abelian Hodge theorem. However, they use a smaller Picard group.
@@
de Cataldo–Zhang also use a smaller Dolbeault side than we do.
@@
Finally, we can state de Cataldo–Zhang’s theorem.
@@
Identifying $\mathcal{S}\simeq\mathcal{M}_{\operatorname{dR}}^{\operatorname{reg}}$ as in Lemma 6.23 and using de Cataldo–Zhang’s Hitchin morphism $\mathcal{M}_{\operatorname{dR}}\xrightarrow{h_{\operatorname{dCZ}}}\mathcal{A}^{\operatorname{dCZ}},$ we can view $\mathcal{S}$ as a $\mathcal{A}^{\operatorname{dCZ}}$-stack.
Theorem 6.38 (de Cataldo–Zhang [15]).
@@
$\mathcal{S}\times^{\widetilde{\mathcal{P}^{\operatorname{dCZ}}}}\widetilde{\mathcal{M}_{\operatorname{Dol}}^{\operatorname{dCZ}}}\to\mathcal{M}_{\operatorname{dR}},$
@@ 6.5. Residues and parabolic structures
6.5.1. Deducing de Cataldo–Zhang’s theorem